OpenStudy (anonymous):

f(x) = x^2 + 3; g(x) = Square root of x-2 Find f(g(x)). please help

OpenStudy (anonymous):

\[f(x)=x^2+3\] \[g(x)=\sqrt{x-2}\] plug in g(x) in f(x) \[f(g(x))=fog(x)=(g(x))^2+3\]

OpenStudy (anonymous):

put the value of g(x) in the equation and simplify

OpenStudy (anonymous):

thank you Nishant_Garg for helping me :)

OpenStudy (anonymous):

(g(x))2+3 = (\[\sqrt{x-2}\](x))2+3)

OpenStudy (anonymous):

you have to square g(x) and add 3 to it according to f(x)

OpenStudy (anonymous):

\[(g(x))^2=(\sqrt{x-2})^2\]

OpenStudy (anonymous):

\[(\sqrt{x-2})^{2}\]

OpenStudy (anonymous):

yep! and add 3 to that

OpenStudy (anonymous):

\[(\sqrt{x-2)}^{2}+3\]

OpenStudy (anonymous):

x+1?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you!